OpenStudy (anonymous):

The lateral area of a pyramid with a square base is 240 sq. ft. Its base edges are 12 ft long. Find the slant height of the pyramid.

OpenStudy (anonymous):

(240 - 144)/4 = 24 = (1/2)sh *12 48/12 = sh sh = 4

OpenStudy (jhannybean):

 The equation for square based pyramid is \[\large A = 2bs+b^2\] where b = base, s = slant height in your problem,you're tryingto find the slant height so our equation would be \[\large SL=\frac{A-b^2}{2b}\]A= 240, b= 12 \[\large SL = \frac{240-(12)^2}{(2*12)}\]\[\large SL=\frac{240-144}{24}\]\[\large SL= \frac{96}{24}\]and you can reduce that to get your answer :)